Grade 12Physical Chemistry

the vapour pressure of a pure liquid a at 310c is 120 torr The vapour pressure of this liquid in the solutin with liq b is 72torr Calculate the mole fraction of a in the solution if the mixture obeys raults law

devangi
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
From Rault’s law,
Total Pressure that is
P = Pa + Pb => 120 + 72 => 192 torr.
and
Pa = Xa*P Putting all the values we get,
120 = Xa*192
Xa = 120/192  => 0.628
This is the mole fraction of the gas a.